OpenStudy (anonymous):

a. Convert 450 RPM into radians/second b. Convert 25.5 radians/second into RPM

OpenStudy (anonymous):

How do I tackle these?

OpenStudy (espex):

\[\frac{rad}{sec}=RPM * \frac{2\pi}{60}\]

OpenStudy (anonymous):

Does my calculator need to be in radian mode for both of these questions?

OpenStudy (espex):

No, as far as your calculator is concerned you are just multiplying numbers.

OpenStudy (espex):

radians mode is for working angles.

OpenStudy (anonymous):

Cheers for explaining that, I'll get to working it out then.

OpenStudy (anonymous):

For a) I got 47.12 (2dp)

OpenStudy (espex):

That is a correct answer. :)

OpenStudy (anonymous):

Cool, and for b) I got 243.51 RPM

OpenStudy (espex):

Same answer, nice work.

OpenStudy (anonymous):

Cheers! :)